OpenStudy (anonymous):

a leaky valve on the water overcharges the residents for one gallon of water every 1 1/3 months. the over charged amount w varies directly with time t. find the equation that models this direct variation. how many months it will take for the residence to be over charged 5 gallons of water? A. w= 3/4t; 6 2/3 months B. w=4/3t; 6 2/3 months C. w=3/4t; 3 3/4 D. 4/3t; 3 3/4 months

OpenStudy (anonymous):

1 and 1/3 = 4/3. so w = (4/3)t, where t is months. now use that function to find the second part. 5 = (4/3)t -----> t = 15/4 = 3 3/4 months